OpenStudy (anonymous):

Find the solution of the logarithmic equation log(3x+2)=3

OpenStudy (whpalmer4):

If \[\log_b x = a\]that means that \[b^a = x\] That means \[\log_{10}(3x+2)=3\implies 10^3 = 3x+2\]assuming base 10 for the logarithm.

OpenStudy (whpalmer4):

\[\implies\]means "implies"